Having hard time understanding the process used to solve this problem. I dont have any issues with 2 (x-1)(2x+3) but 3 is throwing me off.... any help would be appreciated...
(x-1)(2x+3)(3x-2)

Answer by smik(40)

Well, it's about the same as you would do it when you're expanding two expressions. When it comes down to expanding 3 expressions, you do two first, and then you multiply the result of that by the one that's left over. In this case, we would multiply (x-1) by (2x+3) to get:


(2x^2 + 3x - 2x - 3) simplified:


(2x^2 + x - 3)


Now we multiply that with (3x-2):


(3x - 2)(2x^2 + x - 3):
3x*2x^2 = 6x^3
3x*x = 3x^2
3x*-3 = -9x
-2*2x^2 = -4x^2
-2*x = -2x
-2*-3 = 6


which gives us as a final answer (when all the like terms are added together):


6x^3 - x^2 - 11x + 6